# Rethinking Knowledge Distillation: A Data Dependent Regulariser With a Negative Asymmetric Payoff

### v1

@@ Abstract
Knowledge distillation is often considered a compression mechanism when judged on the resulting student’s accuracy and loss, yet its functional impact is poorly understood. In this work, we quantify the compression capacity of knowledge distillation and the resulting knowledge transfer from a functional perspective, decoupling compression from architectural reduction, which provides an improved understanding of knowledge distillation. We employ hypothesis testing, controls, and random control distillation to understand knowledge transfer mechanisms across data modalities. To rigorously test the breadth and limits of our analyses, we explore multiple distillation variants and analyse distillation scaling laws across model sizes. Our findings demonstrate that, while there is statistically significant knowledge transfer in some modalities and architectures, the extent of this transfer is less pronounced than anticipated, even under conditions designed to maximise knowledge sharing. Notably, in cases of significant knowledge transfer, we identify a consistent and severe asymmetric transfer of negative knowledge to the student, raising safety concerns in knowledge distillation applications. Across 12 experimental setups, 9 architectures, and 7 datasets, our findings show that knowledge distillation functions less as a compression mechanism and more as a data-dependent regulariser with a negative asymmetric payoff.
@@ 1 Introduction
Large neural networks have achieved remarkable results across domains (Brown et al., 2020; Dosovitskiy et al., 2021; Kirillov et al., 2023), but at significant computational cost. This has motivated techniques that reduce model size while maintaining performance. Knowledge distillation (KD) has emerged as a widely adopted method to compress models by training a student model to mimic a larger teacher (Buciluǎ et al., 2006; Hinton et al., 2015; Gu et al., 2024; Muralidharan et al., 2024). While KD can be applied across architectures and modalities – including in self-distillation regimes where the teacher and student share the same architecture (Allen-Zhu & Li, 2023; Zhang et al., 2019) – the mechanism by which KD improves student performance remains unknown (Busbridge et al., 2025). Recent studies have challenged the assumption that KD works through meaningful knowledge transfer, showing that performance gains have been observed even with randomly initialised teachers (Stanton et al., 2021a) motivating a rigorous examination of KD’s functional impact.
In this work, we move beyond the question of whether knowledge is transferred – we challenge the framing of Knowledge Distillation as a mechanism of knowledge transfer altogether. We argue that the improvements observed do not necessarily arise from meaningful transfer of the teacher’s knowledge, but from a more general, data-dependent regularisation effect disputed in literature (Stanton et al., 2021a; Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020) with a novel identification of a negative asymmetric payoff in KD. To support this claim, we study KD from a functional perspective, and quantify how closely student models replicate the teacher’s output function. We ground our work around two research questions: 1) Does knowledge distillation result in a significantly functionally similar model to the teacher across architectures and data domains against controls? 2) What knowledge, if any, is actually transferred to student models?
We first focus on self-distillation, where the student has the capacity to match the teacher’s functional representation perfectly, ensuring that any observed differences are solely due to the distillation signal. We then verify our findings in the standard distillation setting with smaller student models (Appendix Section E), as well as with different KD variants in Appendix Section C.
@@
Our methodological framework isolates the core mechanics of Knowledge Distillation through: 1) a controlled training setup where all models share initialisation, enabling precise functional comparison; 2) two controls: independent models with the same architecture,initialisation and different data order (SIDDO) as the teacher, and a Random Control Distillation (RCD) where students are trained using uniform noise in place of teacher outputs, all functionally compared to the teacher model used in the standard distillation process; 3) functional similarity metrics including Activation Distance, Rank Disagreement, Prediction Disagreement, JS Divergence and Prediction Agreement.
We conduct experiments across 7 datasets, 3 data modalities (image, audio, and language), and 9 architectures, training over 3,900 models. Our findings show that:
While KD can lead to statistically significant functional similarity between teacher and student, this similarity is often marginal and inconsistent across datasets and modalities.
@@
The most substantial improvements in accuracy and loss frequently arise under Random Control Distillation, challenging the assumption that performance gains reflect successful knowledge transfer.
@@
When knowledge transfer is significant and not marginal, the transferred knowledge has an asymmetric weighting towards the teacher’s incorrect predictions. This asymmetry becomes more pronounced as dependence on the teacher increases.
@@
Our findings compel a re-characterisation of KD, not as a robust knowledge transfer mechanism, but as a data-dependent regulariser with inconsistent and negative asymmetric knowledge-sharing capacity. This perspective raises important safety concerns: when knowledge transfer is significant, KD may amplify incorrect or harmful behaviour encoded in the teacher. We present a concrete case of adversarial transfer facilitated by KD to support this.
Concretely, our contributions are as follows:
Introduce a functional framework to analyse KD beyond accuracy and loss, but as a process where internal knowledge transfer dynamics can be quantitatively measured.
Isolate the contribution of the teacher signal using strong statistical and control-based methodology, something that prior work has not quantitatively disentangled to this level.
@@
Identify and characterise a novel phenomenon across conditions, modalities and architectures: when functional transfer occurs, it disproportionately favours the teacher’s incorrect predictions, revealing a systematic error amplification effect with safety implications.
@@
Demonstrate the diagnostic utility of RCD as a crucial counterfactual, showing it frequently outperforms KD, undermining assumptions about knowledge transfer.
@@
Conduct the largest multimodal functional study of KD to date. Our empirical analysis spans over 3,900 trained models across 9 architectures, 7 datasets, and 3 modalities (vision, audio, and language), establishing the generality and reproducibility of our claims.
@@
Reveal targeted and scalable negative transfer via adversarial and capacity scaling experiments. We show that KD can reliably copy specific erroneous behaviours, and that this error amplification scales with model capacity, underscoring the hidden risks of KD in high-stakes settings.
@@ 2 Related Work
KD transfers behaviour from a teacher (or ensemble) into a student (Buciluǎ et al., 2006; Hinton et al., 2015), with strong empirical results across modalities (Beyer et al., 2022; Jung et al., 2020; Sanh, 2019; Aghli & Ribeiro, 2021; Li et al., 2020; Fang et al., 2021; Wang et al., 2022) and architectures (Touvron et al., 2021; Miles et al., 2024). Yet the role of knowledge transfer is debated (Mason-Williams, 2024; Stanton et al., 2021b; Ojha et al., 2023; Menon et al., 2021). Prior work alternately views KD as a regulariser (Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020) or argues against that view (Shen et al., 2021; Sultan, 2023). In this paper, we advance the discussion surrounding KD as a regulariser with a functional perspective that spans image, audio, and language. We present a control-driven functional protocol that decouples compression from size, measures alignment beyond accuracy, confirming KD acts as a data-dependent regulariser but exposing a new dimension of this regularisation with respect to its systematic negative transfer to the student.
@@
Functional similarity compares models by their outputs rather than only their accuracy (Klabunde et al., 2023). It has been used for unlearning (Golatkar et al., 2021; Chundawat et al., 2023), ensemble dynamics (Fort et al., 2019), and compression/pruning (Mason-Williams & Dahlqvist, 2024; Mason-Williams, 2024). Metrics such as Activation Distance, Prediction Dissimilarity and JS Divergence have been used for functional analysis. Activation Distance represents the distance on the softmax output distribution of two models, enabling functional comparison. In comparison, JS Divergence represents the Jensen-Shannon information-theoretic divergence that employs a weighted average of KL divergence of distributions, giving a directed divergence between non-continuous distributions (Lin, 1991). Prediction Dissimilarity compares the disagreement of label predictions between models, allowing for an enriched perspective on the alignment of the model’s functions (Fort et al., 2019). We employ all of the above to conduct a functional analysis of knowledge transfer in knowledge distillation.
@@ 3 Experimental Setup
We focus primarily on self-distillation, where the student model has the same architecture and initialisation as the teacher. This setting gives the student maximal capacity to recover the teacher’s function, allowing isolation of the effects of the distillation signal itself. This is achieved through architectural and initialisation matching, along with carefully structured control conditions. Our core experimental findings are derived from this controlled self-distillation setup. To verify generality, we replicate our results in the standard KD setting with smaller students (Appendix E) as well as with multiple KD variants in Appendix Section C.
@@
Let denote the teacher model, trained from initialisation . All subsequent models – including students and controls – share the same architecture and initialisation , ensuring they begin from the same point in the loss landscape. Thus any observed differences in functional behaviour arise purely from the training signals (e.g., data order or distillation) rather than confounds from architecture or initialisation. In self-distillation, students start from and are trained to match the finalised teacher with the standard logit-matching objective:
@@
where is the input, is the student model parameters, is the teacher weighting coefficient, is the cross-entropy loss function, is the kullback-leibler divergence loss function, is the ground truth label, is the softmax function parameterised by the temperature , and and are the logits of the student and the teacher, respectively. Unless otherwise stated, we keep all training hyperparameters fixed across conditions: optimiser, learning-rate schedule, batch size, data augmentations/preprocessing, epochs, and evaluation protocol.
To isolate the effect of the teacher signal, we introduce a Random Control Distillation (RCD) setup, analogous to a randomised control trial (Hariton & Locascio, 2018). Here, the student is trained with the same distillation loss (Eq. 1), but the teacher outputs are replaced by samples from a uniform distribution in . This setup is visualised in Figure 2(a).
We vary the distillation coefficient to modulate reliance on the teacher. At 0.1, the teacher signal contributes minimally; at 0.5, there is an equal weighting of label and teacher supervision; at 0.9, training is predominantly guided by the teacher. If KD achieves meaningful knowledge transfer, functional similarity should increase with higher . All experiments use temperature to preserve the original teacher distribution.
For each architecture–dataset pair spanning over different modalities, we train 3 teacher models (seeds 0-2), and 10 student models per distillation setup (KD, RCD, SIDDO; see below) 3 values (seeds 10-19). This results in 73 models per dataset–architecture pair, and a total of 3,942 models across all conditions (Table 1). Results are reported using Standard Error of the Mean (SEM) (Belia et al., 2005), which better reflects estimation uncertainty across independent runs. Our experimental setup is aligned with reproducibility protocols, defined by Mason-Williams & Mason-Williams (2025), using hypothesis testing and levering statistical significance to enhance the robustness and generality of our findings.
| Modality | Datasets | Architectures | ||
| Image | TinyImageNet Le & Yang (2015), CIFAR10 Krizhevsky et al. (2009), SVHN Netzer et al. (2011) |
@@
VGGish Hershey et al. (2017), AST Gong et al. (2021) | ||
| Language | Tiny Shakespeare Blog (2015), Adversarial Tiny Shakespeare (THA) | Nano-GPT, Pico-GPT Karpathy (2022) |
@@ 3.1 Functional Similarity Metrics
We evaluate student–teacher alignment using functional similarity metrics computed on the test set , comparing teacher and comparison model :
Activation distance: distance between softmax outputs of and .
Rank Disagreement: Percentage of disagreement in the sorted output logits.
Prediction Disagreement: Proportion of mismatched top-1 predictions..
Prediction Agreement: Complement of prediction disagreement (used in error analysis).
Jensen-Shannon (JS) Divergence: A weighted average of KL divergence (Lin, 1991) between the softmax outputs of and .
These metrics move beyond accuracy and loss to quantify the extent to which students reproduce the teacher’s output function.

### v2

# A Functional Perspective on Knowledge Distillation in Neural Networks

@@ Abstract
Knowledge distillation is considered a compression mechanism when judged on the resulting student’s accuracy and loss, yet its functional impact is poorly understood. We quantify the compression capacity of knowledge distillation and the resulting knowledge transfer from a functional perspective, decoupling compression from architectural reduction to provide an improved understanding of knowledge distillation. We employ a control-driven experimental protocol with hypothesis testing and random control distillation to isolate and understand knowledge transfer mechanisms across data modalities. To test the breadth and limits of our analyses, we study self-distillation, standard distillation, feature-map matching variants, distillation scaling laws across model sizes, and the impact of temperature on knowledge transfer. We find statistically supported knowledge transfer in some modalities and architectures; however, the extent of this transfer is less pronounced than anticipated, even under conditions that maximise knowledge sharing. Notably, in cases of significant functional transfer, we identify a consistent and severe asymmetric transfer of negative knowledge to the student, raising safety concerns for knowledge distillation. Across 22 experimental setups, 9 architectures, and 7 datasets, our results suggest that knowledge distillation functions less as a robust compression-by-transfer mechanism and more as a data-dependent regulariser whose transfer component is biased towards negative asymmetric transfer.
Keywords: Knowledge Distillation, Functional Similarity, Asymmetric Transfer, Distillation Scaling Laws, Adversarial Transfer
@@ 1 Introduction
Large neural networks have achieved remarkable results across domains (Brown et al., 2020; Dosovitskiy et al., 2021; Kirillov et al., 2023), but at significant computational cost. This has motivated techniques that reduce model size while maintaining performance. Knowledge distillation (KD) has emerged as a widely adopted method to compress models by training a student model to mimic a larger teacher (Buciluǎ et al., 2006; Hinton et al., 2015; Gu et al., 2024; Muralidharan et al., 2024). While KD can be applied across architectures and modalities, including in self-distillation regimes where the teacher and student share the same architecture (Allen-Zhu and Li, 2023; Zhang et al., 2019), the mechanism by which KD improves student performance remains contested (Busbridge et al., 2025). Recent studies have challenged the assumption that KD works through meaningful knowledge transfer, showing that performance gains have been observed even with randomly initialised teachers (Stanton et al., 2021). This motivates a rigorous examination of KD’s functional impact.
In this work, we move beyond the question of whether knowledge is transferred: we re-examine the framing of knowledge distillation as a robust mechanism of knowledge transfer. Our contribution is not only to examine the regularisation view of KD, but to make it testable: we introduce a functional, control-driven evaluation protocol that measures teacher-student functional similarity relative to counterfactual baselines and decomposes transferred agreement into teacher-correct and teacher-incorrect regions. This lets us distinguish apparent KD gains that are consistent with data-dependent regularisation from gains accompanied by measurable teacher-function transfer. Throughout the paper, ‘knowledge transfer’ refers to our operational notion of teacher-function transfer, measured relative to counterfactual controls.
Our study is grounded in two research questions: 1) Does knowledge distillation produce statistically supported teacher-student functional similarity, relative to control conditions, across architectures and data modalities? 2) When functional transfer occurs, is the resulting teacher-student alignment symmetric across teacher-correct and teacher-incorrect behaviour, or is it systematically asymmetric? Across our experiments, we find that improvements under KD do not necessarily arise from meaningful transfer of the teacher’s knowledge, but are often consistent with a substantial data-dependent regularisation effect (Stanton et al., 2021; Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020).Crucially, when functional transfer is statistically supported, we identify a systematic negative asymmetric payoff: KD disproportionately increases agreement on the teacher’s incorrect predictions relative to agreement on its correct predictions.
@@
We first focus on self-distillation, where the student has the capacity to match the teacher’s functional representation perfectly, ensuring that any observed differences are solely due to the distillation signal. We then verify our findings in the standard distillation setting with smaller student models (Section 4.2), as well as with different KD variants (Section 4.3). Our methodological framework isolates the core mechanics of knowledge distillation through: 1) a controlled training setup where all models share initialisations and data orders, enabling precise functional comparison; 2) two counterfactual controls: independent models with the same initialisation seed and different data order (SIDDO) as the teacher (using the same architecture in the self-distillation case), and a Random Control Distillation (RCD) where the distillation target is replaced with a class-uniform distribution, all functionally compared to the teacher model used in the standard distillation process; 3) a broad array of functional similarity metrics including Activation Distance, Rank Disagreement, Prediction Disagreement, JS Divergence and Prediction Agreement to quantify teacher-student functional alignment and functional transfer between each control and the teacher. We conduct experiments across 7 datasets, 3 data modalities (image, audio, and language), and 9 architectures, training over 4,000 models. Our findings show that:
@@
The strongest improvements in accuracy and loss frequently arise under Random Control Distillation, challenging the assumption that KD performance gains necessarily reflect successful knowledge transfer, and supporting a data-dependent regularisation perspective of KD.
KD can lead to statistically supported functional similarity between teacher and student, but this similarity is often marginal and inconsistent across datasets and modalities. Furthermore, increasing the distillation temperature significantly reduces the degree of functional transfer.
@@
When functional transfer is supported and not marginal, the resulting teacher-student alignment is often asymmetric, with disproportionate agreement on the teacher’s incorrect predictions. This asymmetry becomes more pronounced as dependence on the teacher signal increases.
@@
In the regimes we study, our findings motivate a re-characterisation of KD, not as a uniformly robust knowledge transfer mechanism, but as a data-dependent regulariser with inconsistent and negative asymmetric knowledge-sharing capacity. This perspective raises important, practical safety concerns: when functional transfer is significant, KD may amplify incorrect or harmful behaviour encoded in the teacher. We further support this claim via a concrete case of adversarial transfer facilitated by KD. Concretely, our contributions are as follows:
We provide a control-driven agnostic experimental framework that isolates the contribution of the teacher signal and evaluates KD gains against counterfactual baselines rather than performance metrics alone, something that prior work has not quantitatively disentangled to this level.
@@
We identify and characterise a novel phenomenon across conditions, modalities and architectures: when statistically supported functional transfer occurs, KD disproportionately favours the teacher’s incorrect predictions, revealing a systematic error amplification effect with safety implications.
@@
We demonstrate the diagnostic utility of RCD as a crucial counterfactual, showing it frequently outperforms KD, concretely weakening the assumption that KD gains primarily arise from knowledge transfer.
@@
To our knowledge, we conduct the largest functional study of KD to date spanning over 4,000 trained models across 9 architectures, 7 datasets, and 3 modalities (vision, audio, and language), supporting the generality of our empirical observations and reproducibility of our evaluation framework.
@@
We show targeted and scalable negative transfer via adversarial and capacity-scaling experiments, demonstrating that KD can reliably copy specific erroneous behaviours, and that negative asymmetric transfer can increase with student capacity – underscoring the safety implications of KD applications in high-stakes settings.
Our study contributes a timely and nuanced perspective on knowledge distillation by turning competing explanations of KD into an empirically testable question, helping to adjudicate between data-dependent regularisation and reliable knowledge-transfer accounts, while uncovering a systematic error-asymmetric transfer dynamic with important safety implications across modalities.
@@ 2 Related Work
KD transfers behaviour from a teacher (or ensemble) to a student (Buciluǎ et al., 2006; Hinton et al., 2015), with strong empirical results across modalities (Beyer et al., 2022; Jung et al., 2020; Sanh, 2019; Aghli and Ribeiro, 2021; Li et al., 2020; Fang et al., 2021; Wang et al., 2022) and architectures (Touvron et al., 2021; Miles et al., 2024). However, the mechanism behind KD improvements remains debated (Mason-Williams, 2024; Stanton et al., 2021; Ojha et al., 2023; Menon et al., 2021). A line of work interprets KD as a form of data-dependent regularisation (Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020), arguing that limited or no knowledge is transferred under its optimisation objective. Other literature, on the other hand, strongly argues that KD can effectively transfer meaningful knowledge from teacher to student (Shen et al., 2021; Sultan, 2023). In this paper, we advance the discussion of KD as a regulariser by introducing a functional perspective that spans vision, audio, and language. There remains substantial contention in the community about how knowledge distillation fundamentally operates (Busbridge et al., 2025), with misaligned interpretations across data modalities. This motivates our multi-modal, multi-architecture, and multi-dataset study of KD. We present a control-driven functional protocol that decouples compression effects from architectural size reduction, and measures teacher-student alignment for statistical significance beyond only analysing accuracy and loss. Using this framework, we provide evidence that KD often acts as a data-dependent regulariser, while also exposing a new dimension of this regularisation: when substantial and statistically supported knowledge transfer occurs, there is a strong negative asymmetric transfer to the student.
@@
Functional similarity methods compare models by their outputs rather than their accuracy alone (Klabunde et al., 2023). Such analyses have been used in unlearning (Golatkar et al., 2021; Chundawat et al., 2023), understanding ensemble dynamics (Fort et al., 2019), and compression/pruning (Mason-Williams and Dahlqvist, 2024; Mason-Williams, 2024). Metrics such as Activation Distance, Prediction Dissimilarity and JS Divergence have been used for functional analysis. Activation Distance represents the distance between the softmax output distribution of two models, enabling output-space functional comparison. JS Divergence represents the Jensen-Shannon information-theoretic divergence that gives a directed measure of divergence between non-continuous probability distributions by employing a weighted average of the KL divergences from each distribution (Lin, 1991). Prediction Disagreement measures the proportion of inputs for which two models produce different top-1 predictions, providing a complementary view of functional alignment at the label level (Fort et al., 2019). We employ all of the above to conduct a functional analysis of KD.
@@ 3 Experimental Setup
We begin with self-distillation, where the student model has the same architecture and initialisation as the teacher. In theory, this setting gives the student the exact capacity to entirely recover the teacher’s function, allowing isolation of the effects of the distillation signal. Our core experimental findings are derived from this controlled self-distillation setup. We then verify the generality of our findings in standard KD (smaller student; Section 4.2), feature-map matching distillation (Section 4.3) and temperature variants (Section 4.4).
@@
Let denote the trained teacher model, trained from initialisation . All comparative models (including students and controls) share architecture and initialisation , ensuring they begin from the same point in the loss landscape. Thus any observed differences in functional behaviour arise purely from the distillation signal rather than confounds from architecture and initialisation. In self-distillation, students start from and are trained to match the trained teacher with the standard logit-matching objective:
@@
where is the input, is the student model parameters, is the teacher-weighting coefficient, is the cross-entropy loss function, is the kullback-leibler divergence loss function, is the ground-truth label, is the softmax function parameterised by the temperature , and and are the student and teacher logits, respectively. Unless otherwise stated, we keep all training hyperparameters fixed across conditions: optimiser, learning-rate schedule, batch size, data augmentations/preprocessing, number of epochs, and evaluation protocol.
To isolate the effect of the teacher signal, we compare KD against two controls: SIDDO (same initialisation, different data order), using the same architecture and initialisation as the teacher, but trained independently with different training-order randomness; RCD (random control distillation; analogous to a randomised control trial (Hariton and Locascio, 2018)), using the same optimisation as KD (Eq. 1), but where teacher distribution is replaced with a class-uniform target for classes (i.e., uniform teacher logits). The setup is visualised in Figure 2(a). RCD is intended as a counterfactual control that preserves the distillation training form while removing teacher information content.
We vary the distillation coefficient to modulate reliance on the teacher. At 0.1, the teacher signal contributes minimally; at 0.5, there is an equal weighting of label and teacher supervision; at 0.9, training is predominantly guided by the teacher. If KD induces meaningful knowledge transfer, functional similarity should monotonically increase with an increase of the (although optimiser dynamics and task effects may introduce deviations). All experiments use temperature to preserve the original teacher distribution allowing for full representation matching. However, in Section 4.4, we vary temperature to analyse its impact on knowledge transfer. For each architecture–dataset pair across modalities, we train 3 teacher models (seeds 0-2). For each teacher, we then train 10 student models for each distillation setup (KD, RCD, SIDDO) over 3 values (0.1, 0.5, 0.9) for KD and RCD - we use 10 training seeds (10-19). This results in 71 models per architecture-teacher combination, and 2,556 models in the self-distillation experiments alone. For smaller-student KD we train 790 models, while for temperature experiments we train 490 models, for a total of 4,189 models across the paper when including adversarial and feature map matching experiments. A full breakdown of models trained per experimental setup is provided in Appendix Section G. Table 1 provides a list of architectures and datasets used, spanning settings from SVHN to ImageNet. Results are reported using the Standard Error of the Mean (SEM) (Belia et al., 2005), which better reflects estimation uncertainty across independent runs.
| Modality | Datasets | Architectures | ||||||
| Image |
@@
| Language |
Nano-GPT, Pico-GPT (Karpathy, 2022) |
@@ 3.1 Functional Similarity Metrics
We evaluate student–teacher alignment using functional similarity metrics computed on the test set, . This enables an analysis of how similar teacher and student representations are on unseen data. Let denote the teacher and a comparative model (KD student, SIDDO, RCD). For the metrics described below, with the exception of Prediction Agreement, lower () values indicate greater functional similarity. The functional similarity metrics examined in this paper are described as follows:
Activation distance: The mean distance between the softmax output distributions of and on , formally expressed as: .
Jensen-Shannon Divergence (JS Divergence): The mean Jensen–Shannon divergence (weighted average of KL divergence (Lin, 1991)) between the softmax output distributions of and on , formally expressed as:
, where . -
Rank Disagreement: The percentage of disagreement in the hierarchical order of output predictions between and on formally expressed as:
Prediction Disagreement: The proportion of inputs on which and produce different top-1 predictions on , formally expressed as: . We also use Prediction Agreement, the complement of Prediction Disagreement, to analyse prediction transfer, for predication agreement increased values indicate more functional similarity.
These metrics move beyond accuracy and loss by quantifying the extent to which students reproduce the teacher’s output function on a given task, which is essential for understanding student–teacher alignment in practice. In Appendix Section A, we show how the functional analysis used in this paper relates to information-theoretic and geometric perspectives.
3.2 Knowledge Transfer Definitions
In this section, we define what, under the experimental conditions explored in this paper, constitutes meaningful knowledge transfer, how it is expected to manifest in the student model, and how different transfer payoffs should be interpreted. We quantify knowledge transfer through functional similarity between teacher and student, since this directly measures the extent to which the student reproduces the teacher’s output function on unseen data. This enables us to understand how knowledge transfer affects the representation of the student model. Accordingly, when knowledge transfer occurs, we expect increased functional alignment between teacher and student, especially in self-distillation settings where the student has the capacity to match the teacher’s function.
Knowledge transfer:
We operationally define knowledge transfer as a statistically significant increase in teacher–student functional similarity relative to both RCD and SIDDO controls. Concretely, we say that knowledge transfer has occurred when most functional similarity metrics (e.g., Activation Distance, Rank Disagreement, Prediction Disagreement, and JS Divergence) improve significantly when comparing the student to the teacher against these baselines. A decrease in these metrics indicates increased functional similarity between teacher and student under KD. If this criterion is met, the resulting agreement between student and teacher, relative to the baselines, can fall into one of three cases detailed below (see Figure 2 for a diagrammatic example).
Let denote the agreement in top-1 label rate between student and teacher restricted to inputs on which the teacher is correct, and let denote the agreement rate restricted to inputs on which the teacher is incorrect. Let and denote the corresponding agreement rates for a baseline model (e.g., SIDDO or RCD). We then define:
and
These quantities measure how much KD changes teacher–student agreement relative to the control baseline on teacher-correct and teacher-incorrect predictions, respectively. Based on these quantities, transfer can take one of three forms:
Asymmetric payoff:
An asymmetric payoff occurs when the increase in teacher–student prediction agreement, relative to controls, differs between teacher-correct and teacher-incorrect predictions. We report the separate changes in correct-agreement, , and incorrect-agreement, , in order to characterise the structure of transferred agreement.
Negative asymmetric transfer:
We define negative transfer as the regime in which two properties hold simultaneously: (i) functional similarity improves relative to controls, but (ii) the increase in incorrect-agreement exceeds the increase in correct-agreement (negative asymmetry). In other words, the student becomes more functionally similar to the teacher while absorbing proportionally more of the teacher’s mistakes than its correct behaviour.